Suppose a young hairstylist is considering opening her salon. The salon could take in a yearly revenue of $150,000, but she must quit her job as an administrative assistant which paid $45,000 a year. She also must take over a building that she owns and rents to someone else for $20,000 a year. She could also get a return of $10,500 a year if she did not use her savings. Her expenses at the salon would be $10,000 for materials, $5,000 for utilities, and $40,000 for an assistant.

a.      List her explicit costs and implicit costs. (2 points)

b.     Calculate her yearly accounting and economic profit? Show your work.(2 points)
Answer #1

Answer

a.      List her explicit costs and implicit costs. (2 points)

explicit costs are the cost paid directly and it is the expanses of the business

expenses at the salon would be $10,000 for materials, $5,000 for utilities, and $40,000 for an assistant

Explicit cost =10000+5000+40000=55000

Implicit cost is costs not paid but these the forgone benefits of the activity

Implicit cost =salary as administrative assitant+building rent + earning on savings

=45000+20000+10500

=75500

b.     Calculate her yearly accounting and economic profit? Show your work.(2 points)

Accounting profit =Revenue - Explicit costs

=150000-55000

=95000

Economic profit =accounting profit -implicit costs

=95000-75500

=19500

The economic profit is $19500